# LORE: A Large Generative Model for Search Relevance

### v1

@@ 1. Path Construction
(2) Mapping Needs to Attributes via Reasoning. While some constraints are explicit, some often lack direct rules, such as constraints on material, style, and color. The model must rely on semantic understanding and reasoning, combined with e-commerce domain knowledge and context, to flexibly align natural language expressions with the structured attribute space.
@@ 3.1 Overview
However, simple SFT was soon proven to suffer from out-of-distribution performance degradation. Researchers discovered that leveraging the model’s CoT capabilities—prompting the model to reason before outputting answers—could effectively enhance model robustness and performance on challenging samples. Due to constraints in computational resources and application scenarios, models with relatively modest parameter sizes are typically selected for training, which implies insufficient reasoning capabilities and instruction-following abilities under the prompting-only paradigm. Therefore, the current mainstream approach is to leverage powerful teacher models (such as GPT-4 and DeepSeek-R1(deepseek)) to synthesize CoT data, and then distill the reasoning capabilities into relatively smaller models through SFT. Motivated by the remarkable achievements of reinforcement learning in LLM training, some works further incorporate an RL stage following CoT distillation, employing either offline RL methods such as Direct Preference Optimization(DPO)(DPO) and KTO(KTO), or online reinforcement learning methods such as Group Relative Policy Optimization(GRPO)(GRPO).
@@ 3.2.1 Feature Construction
First, to address the need for visual attributes, such as stylistic details, cuts and patterns, we incorporate the item’s Main Product Image. This visual input is essential for “image-heavy” queries where textual descriptions alone are insufficient. Additionally, to enrich the item-side attribute features, we introduced category-property-value (CPV) triples and Stock-keeping-unit(SKU). Considering that CPV information contains some redundant items that are irrelevant for relevance judgment, we introduce a pipeline combining prior knowledge with LLMs for filtering. Specifically, CPV information is presented in key-value pairs, and there exists key co-occurrence across similar products. For example, both ”eye cream” and ”face cream” products typically contain the key ”brand”. Therefore, we conducted statistical counting based on a large volume of product data to obtain medium-to-high frequency keys. Subsequently, we prompted a powerful LLM Qwen3-235B-Instruct with the relevance task definition, the considered attributes and each key, asking the model to judge whether each attribute is beneficial for the relevance task. To improve judgment accuracy, we performed eight sampling iterations on the model outputs and retained only those attributes where the self-consistency result was ”true”, which can be described as follows:
@@
However, even after cleaning, CPV data cannot completely avoid noise introduced by item information quality issues. Furthermore, item SKU features exhibit significant redundancy in the feature space due to their inherent attribute characteristics. Therefore, we extract a portion of the data and partition it into training and test sets to conduct feature effectiveness validation experiments. The experiments use product titles as the baseline and progressively introduce other features for incremental analysis. As shown in Table 2, the introduction of both CPV and SKU features improves model performance, indicating that as long as the input information possesses relevance and stability, it can still yield performance gains for the model even in the presence of partial redundancy and noise.
@@
Besides, we introduced item main images to provide visual features that cannot be fully covered by text alone. We exclude SKU images because the main image already captures the majority of visual information. SKU images typically represent minor variations (e.g., style, color) that are adequately conveyed through SKU textual features. Including them would increase modeling complexity without meaningful information gain. Finally, we obtained a rich product information repository, which will be used in all subsequent training.
@@ 3.2.3 Prompt Optimization
In domain-specific SFT, we typically fix the common task definition and requirements as a template, and fill in the query and item information from each data sample to generate the prompt. To effectively activate the core capabilities of the base model, it is necessary to incorporate key information such as relevance task definitions, judgment rules, and output formats into the prompt. Additionally, to guide the model in effectively following our proposed two-stage discrimination framework, we incorporated structured instructional information into the prompt: first requiring the model to understand the query’s intent and attribute requirements; then extracting relevant attributes from item information based on these requirements and making judgments according to the rules. For reasoning tasks, the more detailed and specific the instructions in the prompt, the easier it is for the model to derive the correct answer. However, this pattern does not necessarily hold in Supervised Fine-Tuning (SFT) scenarios, as the model primarily acquires capabilities by learning from response data rather than solely relying on instructions. To obtain the optimal prompt configuration, we explored three versions of prompt designs. The long version includes relevance task definitions, judgment requirements, and detailed discrimination rules, totaling over 7,000 tokens. The mid version streamlines the long version by removing detailed rules and retaining only the most essential instructions, with a length of approximately 800 tokens. The short version is the most concise, preserving only the basic role definition, such as: ”You are an e-commerce relevance assessment expert. Please score the query and item.”
Table 3: Experimental results of SFT across different prompt lengths.
Models pass@1
@@
long prompt 0.866
We validated using the training and test sets mentioned in Section 3.2.1, with results shown in Table 3. The performance differences among the three prompt versions after training are minimal (within 1%), indicating that for single-task SFT scenarios, model capability improvement primarily stems from learning the ground truth, with relatively limited impact from the prompt. Specifically, the short prompt achieves the worst results as expected, due to its lack of necessary task information, causing the model to merely fit the training data distribution. Surprisingly, the long prompt, which contains rich rule details, underperforms compared to the streamlined mid-length prompt (mid prompt). We hypothesize that this is because the excessively long prompt causes attention dispersion in the model, increasing discrimination difficulty. The experimental results demonstrate that a concise prompt containing core information yields optimal performance. Therefore, in subsequent fine-tuning experiments, we uniformly adopt the mid-length prompt configuration.
@@ 3.3.1 Data construction
This phase represents the model cold start stage, which aims to enhance the model’s fundamental discriminative capability. To achieve this objective, it is necessary to construct a training dataset featuring two core characteristics: comprehensiveness and low noise. Regarding the former, we need a scientific sampling strategy to adequately cover diverse e-commerce data distributions, thereby avoiding oversampling or undersampling phenomena for specific data subsets. Concerning the latter, given that the final data annotation relies on manual completion and errors and omissions during the manual annotation process are difficult to completely avoid, empirical testing demonstrates that the annotation accuracy only reaches 95%. In light of this, we design a data cleaning pipeline to systematically reduce data noise and improve dataset quality.
@@
Data sampling. Based on Taobao search logs, we first categorize the data into industries such as consumer electronics, apparel, etc., and conduct stratified sampling according to the proportion of each industry. The purpose of this step is to ensure that the training data cover attributes and features across all industries. Next, we divide the data for each industry into high/medium/low frequency categories based on query frequency and conduct sampling according to their respective proportions, ensuring that the dataset encompasses both common high-frequency queries and long-tail difficult samples. Subsequently, the data is assigned to specialized annotators for labeling.
Data cleaning. The errors in manual annotation primarily stem from two aspects: first, annotators’personal oversights lead to the omission of certain detailed information, a phenomenon particularly pronounced when processing long multi-attribute queries; second, the annotation rules themselves exhibit ambiguity, and as concrete expressions of users’ subjective preferences, existing rules struggle to exhaustively cover all edge cases. To address these two types of issues, we leverages powerful language models for data filtering: (1) the attribute matching model tends to execute strict matching strategies, demonstrating superior fine-grained feature capture capabilities compared to manual annotation in multi-attribute scenarios; (2) by injecting complete annotation rules into prompts and guiding the model to perform reasoning and judgment following instructions, corner cases beyond the scope of rule coverage are effectively identified and filtered out.
Specifically, First, we prompt a powerful text model, Qwen3-235B-Instruct, with detailed definitions of relevance tasks and rules, instructing the model to make 8 separate judgments on each pair $Q_{i},I_{i}>$. Subsequently, we prompt a powerful VLM, Qwen2.5-72B-VL with the same settings and additionally incorporated product main images to obtain eight judgment results for each $Q_{i},I_{i}>$ pair. We merged the results from both inference rounds, totaling 16 judgments, and filtered out samples where none of the judgments agreed with the manual annotations, as these are highly likely to be incorrectly labeled samples or Extremely difficult samples. This process can be represented as follows:
@@
The rationale for employing two models for filtering is to leverage the powerful long-context comprehension capability of the text model and the multimodal understanding capability of the VLM, thereby reducing both false positives and false negatives in the sample filtering process. Ultimately, this study submits the aforementioned filtered suspicious samples to a rework process, where annotators perform categorical processing: for pure annotation errors, the annotation results are directly corrected; for annotation discrepancies caused by rule ambiguity, new annotation guidelines are formulated through the induction and summarization of similar cases, thereby achieving iterative optimization of the annotation rule system. Table 4 presents the results after data cleaning. The cleaned samples involved modifications to 6.1% of the labels, and following this quality inspection, the data quality improved from the original 95% to 99%.
@@ 3.3.2 Multi-dimensional Chain-of-Thought Synthesis
To enable the model to possess comprehensive and in-depth reasoning capabilities, we proposes a progressive CoT synthesis pipeline as shown in Fig 8. This pipeline adopts a stage-wise construction strategy, sequentially encompassing three progressive stages:Knowledge Injection & Reasoning, Multimodal Comprehension, and Rule Awareness. Knowledge Injection & Reasoning completes the Path Construction stage, which parses query intent into attributes comparable with items. Multimodal Comprehension and Rule Awareness then complete the Path Following stage, which evaluates the degree to which items satisfy the query based on item information and discrimination rules.
(1) Knowledge Injection and Reasoning Foundation: Domain knowledge constitutes the cognitive foundation for understanding user queries and item attributes, necessitating injection at the outset. Building upon this knowledge base, the model can achieve deep understanding of query intent, accurately infer users’ core requirements, and accordingly extract corresponding attribute features from item information.
(2) Multimodal Information extraction: the visual presentation of items often contains critical features that are difficult to fully express through text (such as style details, material texture). In light of this, this stage further introduces item image information, synthesizing attribute extraction chains of thought that can integrate both textual and visual modalities. Through cross-modal information complementarity and verification, the model can construct more comprehensive and accurate item representations.
(3) Rule-Oriented Discrimination: Following the completion of query intent recognition and item attribute extraction, attribute matching and discrimination still depend on objective rules. This stage synthesizes rule-aware reasoning chains, enabling the model to explicitly comprehend, adhere to, and apply these discrimination rules throughout the reasoning process.
Step 1: Knowledge and Reasoning. In Section 2.3, we categorized knowledge into general world knowledge and e-commerce knowledge. For the former, existing LLMs have already accumulated considerable knowledge through large-scale pre-training, enabling them to understand common concepts and entity relationships. However, e-commerce scenarios contain numerous phrases and expressions that are difficult to comprehend based solely on their independent semantic information.
@@
Such issues can be primarily categorized into the following two dimensions: (1) Long-tail Query: These queries explicitly refer to extremely niche specific items, yet existing models lack the corresponding prior knowledge reserves, making it difficult to establish effective semantic mapping relationships. For instance, the query is ”strange things are also loved by someone” and the item is a ”novel and interesting toys,” the query actually refers to the title of a book. Without this prior knowledge, the model is challenging to make correct relevance judgments based purely on literal semantics. Similar cases include brand slang, product nicknames, marketing terminology, and other e-commerce-specific expressions; (2) Semantically Ambiguous Items: Certain items fall outside the model’s knowledge boundaries due to their excessively niche nature. For instance, the term ”lying gate” cannot be accurately identified and categorized based solely on its textual title information without adequate contextual understanding of the product attributes. These contents require supplementary domain-specific e-commerce knowledge for accurate interpretation.
@@
Upon completion of knowledge augmentation, the model has acquired the fundamental capability to perform deep-level reasoning on query intent and attributes. This study employs the Qwen3-235B-Instruct model to generate reasoning paths. Through carefully designed prompt templates, the model is guided to execute multi-level reasoning analysis: the model first conducts semantic understanding of the query by integrating its inherent knowledge with e-commerce domain knowledge, inferring users’ true intent and extracting specific attribute features; subsequently, it performs systematic analysis of item information, deconstructing the item’s concrete attributes and establishing comparative mappings with the query. This process generates reasoning chains $CoT_{step1}$ that comprehensively understand both queries and items as follows:
@@
Step 2: Multimodal Comprehension. The $CoT_{step1}$ generated above only encompasses the reasoning process for the text modality and lacks integration of image modality. Based on the following two considerations, we did not directly employ VLM for multimodal modeling: (1) Limited reasoning capability: Existing VLMs generally exhibit weaker reasoning capabilities compared to LLMs of equivalent parameter scale, a phenomenon that has been validated in multiple benchmark tests and our experiments. (2) Modality dependence bias: Direct fine-tuning of VLMs tends to cause the model to exhibit a ”text shortcut” phenomenon, i.e., over-reliance on textual information while neglecting image features. The underlying cause lies in the inherent information density disparity between text and image modalities.
Therefore, we designed a caption-mediated two-stage modeling framework: in the first stage, we extract image semantics through VLM and generate textual descriptions; in the second stage, these descriptions are incorporated into the prompt as additional context for fine-tuning. However, naive caption generation methods suffer from inherent flaws: due to the complex information in images and lack of task guidance, VLMs struggle to a priori locate key information relevant to reasoning, leading to the omission of important visual cues, which cannot be fully resolved through simple prompt engineering. Therefore, we designed a multimodal CoT generation scheme guided by relevance judgment cues.
Firstly, We prompt a Qwen2.5VL-72B model to generate a basic caption $C_{1}$ for the image as follows:
@@
Secondly, We integrate the original input $Caption_{naive}$, the $CoT_{step1}$ obtained from the previous section, and other contextual information into a structured prompt, and input it to Qwen2.5-VL-72B to regenerate the caption. This design enables the model to extract task-relevant visual information from the image in a targeted manner under the guidance of explicit reasoning key points, thereby generating more precise and task-oriented descriptions.
@@
Step3: Rule-Aware. The $CoT_{step2}$ obtained in the previous section contains the understanding of both query and Item, essentially completing the step of ”query reaching the product.” However, as described in Section 2.3, the degree to which a product satisfies a query is determined by objective and concrete judgment rules, which requires the model to possess rule-aware discrimination capability. Directly writing the complete rules into the prompt would result in excessively long context, increasing the difficulty of instruction-following for the model. Therefore, we integrated the rules into the CoT, enabling the model to learn this reasoning approach and thereby acquire an understanding of the rules. Specifically, we partition the rule system by industry to obtain industry-specific rule sets R. For each training sample, we construct a prompt comprising: (1) the rule set R, (2) the $context_{i}$ , (3) the ground-truth $L_{i}$, and (4) the previously generated $CoT_{step2}$. This prompt guides the Qwen3-235B-Instruct model to reverse-engineer a rule-based $CoT_{step2}$ for relevance judgment by integrating the complete rule framework, human annotations, and deep semantic comprehension of both query and item contexts.

### v2

@@ 1. Path Construction
(2) Mapping Needs to Attributes via Reasoning. While some constraints are explicit, many others—particularly those related to subjective attributes like material, style, and color—are often implicit and lack direct matching rules. Therefore, the model must integrate semantic understanding and logical reasoning with e-commerce domain knowledge and contextual cues. This allows it to flexibly align natural language expressions with the structured attribute space.
@@ 3.1 Overview
However, simple SFT was soon proven to suffer from out-of-distribution performance degradation. Researchers discovered that leveraging the model’s CoT capabilities—prompting the model to reason before outputting answers—could effectively enhance model robustness and performance on challenging samples. Due to constraints in computational resources and application scenarios, models with relatively modest parameter sizes are typically selected for training, which implies insufficient reasoning capabilities and instruction-following abilities under the prompting-only paradigm. Therefore, the current mainstream approach is to leverage powerful teacher models (such as GPT-4 and DeepSeek-R1(deepseek)) to synthesize CoT data, and then distill the reasoning capabilities into relatively smaller models through SFT. Motivated by the remarkable achievements of reinforcement learning in LLM training, some works further incorporate an RL stage following CoT distillation, employing either offline RL methods such as Direct Preference Optimization(DPO)(DPO) and KTO(KTO), or online RL methods such as Group Relative Policy Optimization(GRPO)(GRPO).
@@ 3.2.1 Feature Construction
First, to address the need for visual attributes, such as stylistic details, cuts and patterns, we incorporate the item’s Main Product Image. This visual input is essential for “image-heavy” queries where textual descriptions alone are insufficient. For textual attributes, we begin by introducing Category-Property-Value (CPV) triples, which provides a rich, semi-structured summary of an item’s core characteristics. Considering that CPV information contains some redundant items that are irrelevant for relevance judgment, we introduce a pipeline combining prior knowledge with LLMs for filtering. Specifically, CPV information is presented in key-value pairs, and there exists key co-occurrence across similar products. For example, both ”eye cream” and ”face cream” products typically contain the key ”brand”. Therefore, we conducted statistical counting based on a large volume of product data to obtain medium-to-high frequency keys. Subsequently, we prompted a powerful LLM (Qwen3-235B-Instruct) with task definition, the considered attributes and each key, asking the model to judge whether each attribute is beneficial for the relevance task. To improve judgment accuracy, we performed eight sampling iterations on the model outputs and retained only those attributes where the self-consistency result was ”true”, which can be described as follows:
@@
Moreover, CPV information alone is insufficient, as it often fails to capture the fine-grained, purchasable variations—such as different colors, sizes, or package contents. To bridge this gap, we incorporate Stock Keeping Unit (SKU) information. SKU provide details for these specific variants, enabling our model to make precise assessments.
Even after the deduplication process, the resulting CPV data can still be noisy due to the inherent quality issues of the original item information. Meanwhile, SKU information often suffers from textual repetition and redundancy. To validate the effectiveness of the features we introduced, we designed a set of incremental analysis experiments. These experiments use the product title as a performance baseline and progressively add other features to evaluate their individual impact on model performance. As shown in Table 2, introducing both CPV and SKU features on top of the title improves model performance. This result indicates that even in the presence of partial redundancy and noise, input information can still yield significant performance gains as long as it contains sufficiently relevant and stable signals.
@@ 3.2.3 Prompt Optimization
In domain-specific SFT, prompts are typically constructed using a fixed template containing task definitions and requirements, populated with query and item data from individual samples. To fully activate the base model’s core capabilities, it is crucial to embed key elements such as relevance definitions, judgment rules, and output formats. Furthermore, to align the model with our proposed two-stage discrimination framework, we integrated structured instructions requiring the model to first analyze the query’s intent and attribute requirements, and subsequently extract relevant item attributes to render a judgment based on established rules.
While detailed instructions generally enhance performance in reasoning tasks, this correlation does not strictly apply to SFT scenarios, where models acquire capabilities primarily through learning from response data rather than relying solely on prompt instructions. To identify the optimal configuration, we evaluated three prompt designs: a long version (about 7,000 tokens) containing exhaustive rules; a mid version (about 800 tokens) retaining only essential instructions; and a short version consisting solely of a basic role definition (e.g., “You are an e-commerce relevance assessment expert…”).
Table 3: Experimental results of SFT across different prompts.
Prompt pass@1
@@
long prompt 0.868
We validated using the training and test sets mentioned in Section 3.2.2, with results shown in Table 3. Model performance showed minimal variance across the three prompt versions (within 1%), indicating that in SFT scenarios, model improvements stem primarily from learning ground truth responses rather than relying on prompt engineering. Specifically, the short prompt yielded the poorest results as expected; its lack of essential task information likely forced the model to merely fit the data distribution without true understanding. Surprisingly, the long prompt, despite its rich detail, underperformed the streamlined mid prompt. We hypothesize that excessive length causes attention dispersion, thereby complicating the discrimination process. The results demonstrate that a concise prompt containing only core information yields optimal performance. Consequently, we adopted the mid-length prompt configuration for all subsequent fine-tuning experiments.
@@ 3.3.1 Data construction
This phase constitutes the model cold start stage, designed to establish the model’s fundamental discriminative capabilities. Achieving this requires a training dataset characterized by two core attributes: comprehensiveness and low noise. To ensure comprehensiveness, we employ a scientific sampling strategy that adequately covers diverse e-commerce data distributions, thereby mitigating potential oversampling or undersampling bias. Addressing noise is equally critical; since manual annotation inevitably introduces errors—with empirical tests showing an accuracy of only 95%—we implemented a robust data cleaning pipeline to systematically reduce noise and enhance overall dataset quality.
@@
Data sampling. Based on Taobao search logs, we first categorize the data by industry-level such as consumer electronics, apparel, etc., and conduct stratified sampling according to the proportion of each industry. The purpose of this step is to ensure that the training data cover attributes and features across all industries. Next, we divide the data for each industry into high/medium/low frequency categories based on query and conduct sampling according to their respective proportions, ensuring that the dataset encompasses both common high-frequency queries and long-tail difficult samples. Subsequently, the data is assigned to specialized annotators for labeling.
Data cleaning. Errors in manual annotation primarily stem from two sources. First, human oversight often leads to the omission of details, a tendency particularly pronounced when processing long, multi-attribute queries. Second, rule ambiguity poses a challenge; as manifestations of subjective user preferences, existing rules struggle to exhaustively cover all edge cases. To address these issues, we leverage powerful language models for data filtering: (1) The attribute matching model executes strict matching strategies, demonstrating superior fine-grained feature capture compared to manual annotation; (2) By injecting comprehensive rules into prompts to guide reasoning, the model could effectively identify and filter out corner cases that fall beyond the scope of standard rules.
Specifically, we prompt a powerful text model, Qwen3-235B-Instruct, with detailed definitions of relevance tasks and rules, instructing the model to make 8 separate judgments on each pair $Q_{i},I_{i}>$. Subsequently, we prompt a powerful VLM, Qwen2.5-72B-VL with the same settings and additionally incorporated product main images to obtain eight judgment results for each $Q_{i},I_{i}>$ pair. We merged the results from both inference rounds, totaling 16 judgments, and filtered out samples where none of the judgments agreed with the manual annotations, as these are highly likely to be incorrectly labeled samples or Extremely difficult samples. This process can be represented as follows:
@@
The rationale for employing a dual-model approach is to synergize the text model’s robust long-context comprehension with the VLM’s multimodal capabilities, thereby minimizing both false positives and false negatives during filtering. Subsequently, suspicious samples identified by this process undergo manual review and are categorized for remediation: clear annotation errors are corrected directly, while discrepancies arising from rule ambiguity trigger the formulation of new guidelines based on case analysis, facilitating the iterative refinement of the annotation rule system. Table 4 presents the post-cleaning results. The process resulted in label modifications for 6.1% of the samples, elevating overall data quality from 95% to 99%.
@@ 3.3.2 Multi-dimensional Chain-of-Thought Synthesis
To enable the model to possess comprehensive and in-depth reasoning capabilities, we proposes a progressive CoT synthesis pipeline as shown in Fig 8. This pipeline adopts a stage-wise construction strategy, sequentially encompassing three progressive stages: (1)Knowledge Injection & Reasoning, (2)Multimodal Comprehension, and (3)Rule Aware. Knowledge Injection & Reasoning completes the Path Construction stage, which parses query intent into attributes comparable with items. Multimodal Comprehension and Rule Awareness then complete the Path Following stage, which evaluates the degree to which items satisfy the query based on item information and discrimination rules.
Step 1: Knowledge Injection & Reasoning. knowledge serves as the cognitive foundation for interpreting user queries and item attributes, necessitating its injection at the initial stage. By leveraging this knowledge base, the model can achieve a profound understanding of query intent, accurately infer core user requirements, and extract relevant attribute features from item information. As outlined in Section 2.3, we categorize knowledge into general world knowledge and e-commerce knowledge. Regarding the former, existing LLMs have acquired substantial general knowledge through large-scale pre-training, enabling them to grasp common concepts and entity relationships. However, e-commerce scenarios frequently involve specialized terminology and phrasing whose meanings cannot be fully comprehended based solely on literal semantics.
@@
Such challenges can be primarily categorized into the following two dimensions: (1) Long-tail Queries: These queries explicitly refer to highly niche items. However, existing models often lack the necessary prior knowledge, making it difficult to establish effective semantic mappings. For instance, consider the query ”strange things are also loved by someone.” While this refers to a specific book title, a model relying solely on literal semantics might incorrectly associate it with ”novel and interesting toys.” Without specific knowledge—such as brand slang, product nicknames, or marketing terminology—it is difficult for the model to make accurate relevance judgments. (2) Semantically Ambiguous Items: Certain items fall outside the model’s knowledge scope due to their extremely niche nature. For example, a term like ”lying gate” cannot be accurately identified or categorized based on textual title information alone; it requires a deeper contextual understanding of product attributes. Ultimately, both cases necessitate supplementary domain-specific e-commerce knowledge for accurate interpretation.
@@
Upon completion of knowledge augmentation, the model has acquired the fundamental capability to perform deep-level reasoning on query intent and attributes. We employs the Qwen3-235B-Instruct model to generate reasoning paths. Through carefully designed prompt templates, the model is guided to execute multi-level reasoning analysis: first conducts semantic understanding of the query by integrating its inherent knowledge with e-commerce domain knowledge, inferring users’ true intent and extracting specific attribute features; subsequently, it performs systematic analysis of item information, deconstructing the item’s concrete attributes and establishing comparative mappings with the query. This process generates reasoning chains $CoT_{step1}$ that comprehensively understand both queries and items as follows:
@@
Step 2: Multimodal Comprehension. The visual presentation of items often contains critical features that are difficult to fully express through text (such as style details, material texture). In light of this, this stage further introduces item image information, synthesizing attribute extraction chains of thought that can integrate both textual and visual modalities. Based on the following two considerations, we did not directly employ VLM for multimodal modeling: (1) Limited reasoning capability: Existing VLMs generally exhibit weaker reasoning capabilities compared to LLMs of equivalent parameter scale, a phenomenon that has been validated in multiple benchmark tests and our experiments. (2) Modality dependence bias: Direct fine-tuning of VLMs tends to cause the model to exhibit a ”text shortcut” phenomenon, i.e., over-reliance on textual information while neglecting image features. The underlying cause lies in the inherent information density disparity between text and image modalities.
Therefore, we designed a caption-mediated two-stage modeling framework: in the first stage, we extract image semantics through VLM and generate textual descriptions; in the second stage, these descriptions are incorporated into the prompt as additional context for fine-tuning. However, conventional caption generation methods face inherent limitations. Given the complexity of visual content and the absence of task-specific guidance, VLMs struggle to effectively identify key information relevant to reasoning. This leads to the omission of critical visual cues—a challenge that cannot be fully addressed through simple prompt engineering. Consequently, we propose a multimodal CoT generation framework guided by relevance judgment.
Firstly, We prompt a Qwen2.5VL-72B model to generate a basic caption $C_{naive}$ for the image as follows:
@@
Secondly, We integrate the original input $C_{naive}$, the $CoT_{step1}$ obtained from the previous section, and other contextual information into a structured prompt, and input it to Qwen2.5-VL-72B to regenerate the caption. This design enables the model to extract task-relevant visual information from the image in a targeted manner under the guidance of explicit reasoning key points, thereby generating more precise and task-oriented descriptions.
@@
Step3: Rule-Aware. While the previous $CoT_{step2}$ effectively bridges the semantic gap, it lacks the objective verification mandated by Section 2.3. To address this, we synthesize rule-aware reasoning chains to enforce rigorous attribute matching. Rather than explicitly listing comprehensive rules—which would overwhelm the context window and hinder instruction following, we embed these constraints directly into the reasoning steps. This allows the model to internalize the judgment criteria implicitly through demonstration. Specifically, we partition the rule system by industry to obtain industry-specific rule sets R. For each training sample, we construct a prompt comprising: (1) the rule set R, (2) the $context_{i}$ , (3) the ground-truth $L_{i}$, and (4) the previously generated $CoT_{step2}$. This prompt guides the Qwen3-235B-Instruct model to reverse-engineer a rule-based $CoT_{step2}$ for relevance judgment by integrating the complete rule framework, human annotations, and deep semantic comprehension of both query and item contexts.